# Uncertainty Quantification for Prior-Data Fitted Networks using Martingale Posteriors

### v1

@@ 1 Introduction
Prior-data fitted networks (PFNs) are foundation models [11, 19] that allow for in-context learning, i.e., the ability to learn at inference time without any parameter updates [9]. TabPFN, a transformer pre-trained on synthetic data for in-context learning on tabular data sets, has recently attracted a lot of interest. TabPFN [11, 12] and extensions such as TuneTables [7] and LocalPFN [24] have been shown to achieve state-of-the-art performance on tabular benchmarks by pre-training on purely synthetic data. Since PFNs and extensions learn in-context, there is no need for further model (fine-)tuning on the inference task.
Recent extensions of PFNs allow their applicability to large data sets [7], the use of PFN “priors” for latent variable models [22], and simultaneously minimizing bias and variance to improve their performance [17]. PFNs are also related to simulation-based inference and amortized inference but have slightly different goals and do not amortize across a single but multiple data sets [22]. While introduced as a Bayesian method and approximation to the posterior predictive, PFNs can also be interpreted as pre-tuned untrained predictors [20]. This also relates to the question of what uncertainty PFN models can provide.
@@
    We introduce a new extension of the MP framework of [8] for inference of predictive quantities conditional on a specific feature value $x$.
@@ 2.1 Prior-data fitted networks
which quantifies the likelihood of observing label $y$ given that the feature is $x$ and $\mathcal{D}_{n}$ has been observed. The PPD is a Bayesian concept and implicitly involves a prior over the distributions $P$ that could have generated the data. To approximate the PPD with a PFN, a deep neural network—typically a transformer—is pre-trained on simulated data sets with diverse characteristics. After pre-training, the network weights are fixed, and the approximate PPD for a new training set can be computed through a single forward pass without additional training or tuning. The PPD quantifies uncertainty about the label $y$. However, it mixes the aleatoric and epistemic components of uncertainty [14]. From the PPD alone, it is impossible to disentangle these parts. Consequently, PFNs do not provide uncertainty estimates for predictive summaries, such as conditional mean, probabilities, and quantiles.
@@ 2.2 Bayesian inference
[…]
@@ 2.3 Martingale posteriors
Martingale posteriors were recently introduced by [8] as a new method for Bayesian uncertainty quantification. Its core idea is to reverse the direction of the Bayesian inference. In classical Bayesian inference, the posterior is derived from a prior and likelihood, which then implicitly leads to the PPD. MP inference starts from the PPD and leaves the prior $\pi(\theta)$ implicit. An appropriate sampling scheme and Doob’s theorem then allow us to derive posteriors for virtually all quantities of interest (e.g., the conditional mean $\mu(x)$).
@@
$y_{n+1}\sim p(\cdot|y_{1:n}),\quad y_{n+2}\sim p(\cdot|y_{1:(n+1)}),\quad y_{n+3}\sim p(\cdot|y_{1:(n+2)}),\quad\dots,$ (1)
$N$ times, which yields a sample $y_{(n+1):(n+N)}$ drawn from the predictive joint distribution
@@
[…]
where $A$ is any Borel set on the space where the parameter $\theta$ lives. Furthermore, Doob’s theorem [3] implies that $\Pi(\theta|\mathcal{D}_{n})$ coincides with the classical Bayes posterior for the prior $\pi(\theta)$ implicit in the PPD [8].
@@ 3 Efficient martingale posteriors for prior-data fitted networks
Martingale posteriors allow for Bayesian inference directly from the PPD. PFNs approximate the PPD, so using PFNs to construct a martingale posterior seems natural. However, there are two problems. First, modern PFNs are based on transformer architectures that require $\Omega(n^{2})$ operations for a forward pass on a training set size of $n$. Iteratively computing $p(y|y_{1:(n+k)})$ for $k=1,\dots,N$ thus has complexity $\Omega(N^{3})$, which is prohibitive. Second, [6] found that modern transformer-based LLMs substantially deviate from the martingale property
$\mathbb{E}[p(y|y_{1:(n+k)})|y_{1:n}]=p(y|y_{1:n}).$
Also TabPFN fails in this regard. LABEL:{fig:martprop} (experiment description in Section 4.1) shows a QQ-plot of the average CDF resulting from PFN’s PPD at iteration 00 vs. the CDF after $k$ forward iterations of the sampling procedure. The martingale property is satisfied when this distribution remains constant, i.e., all lines lie on the $x=y$ diagonal. However, we see an apparent change in the distribution over the different sampling horizons. This reinforces that PFNs approximate Bayesian inference only in a weak sense, even though they were conceived from Bayesian ideas [11, 18].
Without the martingale property, the MP sampling procedure leads to meaningless results. Instead, we propose to use the PPD implied by the PFN only as a starting point for the sampling scheme. This PPD is then iteratively updated using the nonparametric Gaussian copula approach of [8], which ensures the martingale property. The resulting martingale posterior effectively treats the PFN output as an informed prior, without requiring the PFN to provide coherent posterior updates.
@@ 3.1 Martingale posteriors for conditional inference
We extend the unconditional sampling scheme outlined in the previous section to the conditional inference setting. Fong et al. [8] already proposed one such extension. Their scheme involves forward sampling of the features $x_{(n+1):(n+N)}$. The distribution of the features isn’t of primary interest but significantly complicates the sampling procedure, which the authors resolved through heuristic simplifications. In contrast to [8], we propose to sample only the labels $y_{(n+1):(n+N)}$ conditional on the event that $x_{n+k}=x$, for a fixed value of $x$ and all $k=1,\dots,N$.
@@
[…]
Common examples of the parameter $\theta(x)$ are the conditional mean
[…]
or a conditional $\alpha$-quantile
$\theta(x)=P^{-1}(\alpha|x).$
@@ 3.2 Efficient PPD updates based on the Gaussian copula
Observe that $p_{0}(y)=p(y|x,\mathcal{D}_{n})$ is the PPD approximated by the PFN $g_{\theta}$. However, the following update distributions $p_{1},p_{2},\dots$ are generally intractable. To alleviate this, [8] proposed a computationally efficient, nonparametric method based on Dirichlet Process Mixture Models (DPMMs) and a copula decomposition of the conditional $p_{k}$. Specifically, we set
$P_{k}(y)=(1-\alpha_{n+k})P_{k-1}(y)+\alpha_{n+k}H_{\rho}(P_{k-1}(y),P_{k-1}(y_{n+k})),$ (2)
where $P_{k}$ is the CDF corresponding to $p_{k}$,
$\alpha_{i}=\left(2-\frac{1}{i}\right)\frac{1}{i+1},\quad H_{\rho}(u,v)=\Phi\left(\frac{\Phi^{-1}(u)-\rho\Phi^{-1}(v)}{\sqrt{1-\rho^{2}}}\right),$
and $\Phi$ is the standard normal cumulative distribution function. The method has a hyperparameter $\rho$, corresponding to a bandwidth that smoothes the updates. Fong et al. [8] proposed to tune this by maximizing the likelihood of the updated densities over the observed data. To increase the alignment of the updates with the PFN baseline, we simulate a new sample from the PFN and optimize the bandwidth on the simulated data.
@@ 3.3 Theoretical properties
Despite the simplicity of the updates given in Equation 2, they provide essential theoretical guarantees. The following is a direct consequence of Theorem 3 by [8].
Proposition 3.1.
@@
By de Finetti’s theorem (e.g., Theorem 1.49 in [23]) it then follows that there is a random variable $\Theta$ such that $(z_{1},z_{2},\dots)$ is conditionally iid given $\Theta$. The distribution of $\Theta$ can be interpreted as an implicit prior. In our setting, this prior depends both on the initial PPD implied by the PFN and the Gaussian copula updates specified by Equation 2. In particular, the following result follows immediately from Proposition 3.1 above and Theorem 2.2 of [2].
Proposition 3.2.
@@
Proposition 3.3.
For any $N\geq 0$ and $\epsilon>0$, there is a constant $C\in(0,\infty)$ such that
$\sup_{y}\limsup_{M\to\infty}\Pr\left(|P_{M}(y)-P_{N}(y)|\geq\epsilon\right)\leq 2\exp\left(-\epsilon^{2}(n+N)/8\right).$
The proof, given in Appendix A, is based on the Azuma-Hoeffding concentration inequality for martingales and relies only on the martingale property of the updates. The result quantifies how well the distribution $P_{N}$ approximates $P_{\infty,x}$ after $N$ updates. Setting $N=0$ corresponds to the case where $P_{N}(y)$ equals the initial PPD given only the observed data. In our setting, this is the output from the PFN. If this PPD converges to a fixed distribution as $n\to\infty$, the proposition implies that $P_{\infty,x}(y)$ must have the same deterministic limit. Hence, the martingale posterior contracts at roughly the same speed at which the PPD converges.
Further, the approximation error decays quickly in $N$, so we can expect the sampling scheme to approximate $P_{\infty,x}$ well already after a moderate number of steps. Specifically, the bound implies
$\limsup_{M\to\infty}\mathbb{E}[|P_{M}(y)-P_{N}(y)|]=O\left(\frac{1}{\sqrt{n+N}}\right).$
The posterior of any Bayesian model contracts no faster than $1/\sqrt{n}$, even for simple, unconditional targets. Contraction rates for curve fitting problems are typically much slower [10, Chapter 9]. This suggests that a rather small number of sampling iterations suffices for negligible errors. Indeed, we observe empirically that the procedure stabilizes already after 100-200 iterations, see Figure 2.
@@ 3.4 Computation
In practice, we can only sample finite sequences and replace the MP by its finite approximation. The procedure is summarized in Algorithm 1. The overall run-time complexity of the algorithm is $O(BN)$, where $B$ is the number of replications, and $N$ is the length of one ‘chain’. Based on the considerations in the previous subsection, we recommend generating around $N=200$ forward samples in each replication. The computations are independent across the outer loop iterations ($b=1,\dots,B$) and straightforward to parallelize; the inner loop ($k=1,\dots,N$) must be run sequentially. Except for the initial computation of $\widehat{\operatorname{PPD}}(y\mid x)$, the runtime is independent of the size $n$ of the training data set.
@@
8:     Compute
$\widehat{P}_{N}^{(b)}(y)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\Bigl{\{}y_{n+i}^{(b)}\leq y\Bigr{\}}.$
@@
11:  Compute the estimated Martingale Posterior:
$\widehat{\Pi}\Bigl{(}\theta(x)\in A|\mathcal{D}_{n}\Bigr{)}=\frac{1}{B}\sum_{b=1}^{B}\mathbf{1}\Bigl{\{}\theta^{(b)}(x)\in A\Bigr{\}}.$

### v2

@@ 1 INTRODUCTION
Prior-data fitted networks (PFNs) are foundation models (Hollmann et al.,, 2023; Müller et al.,, 2022) that allow for in-context learning, i.e., the ability to learn at inference time without any parameter updates (Garg et al.,, 2022). TabPFN, a transformer pre-trained on synthetic data for in-context learning on tabular datasets, has recently attracted a lot of interest. TabPFN (Hollmann et al.,, 2023, 2025) and extensions such as TuneTables (Feuer et al.,, 2024) and LocalPFN (Thomas et al.,, 2024) have been shown to achieve state-of-the-art performance on tabular benchmarks by pre-training on purely synthetic data. Since PFNs and extensions learn in-context, there is no need for further model (fine-)tuning on the inference task.
Recent extensions of PFNs allow their applicability to large datasets (Feuer et al.,, 2024), the use of PFN “priors” for latent variable models (Reuter et al.,, 2025), and simultaneously minimizing bias and variance to improve their performance (Liu and Ye,, 2025). PFNs are also related to simulation-based inference and amortized inference but have slightly different goals and do not amortize across a single but multiple datasets (Reuter et al.,, 2025). While introduced as a Bayesian method and approximation to the posterior predictive, PFNs can also be interpreted as pre-tuned untrained predictors (Nagler,, 2023). This also relates to the question of what uncertainty PFN models can provide.
@@
    We introduce a new extension of the MP framework of Fong et al., (2023) for inference of predictive quantities conditional on a specific feature value $x$.
@@
    We extend existing learning rate schedules and discuss the role of contraction rates in light of inconsistent PFN estimators.
  4. 4.
@@
Refer to caption Refer to caption Refer to caption
Figure 1: Visualization of the diffusion process data (left) together with the PFN predicted density (center) and the estimated epistemic uncertainty (right) via our martingale posterior approach.
@@ 2.1 Prior-Data Fitted Networks
which quantifies the likelihood of observing label $y$ given that the feature is $x$ and $\mathcal{D}_{n}$ has been observed. The PPD is a Bayesian concept and implicitly involves a prior over the distributions $P$ that could have generated the data. To approximate the PPD with a PFN, a deep neural network—typically a transformer—is pre-trained on simulated datasets with diverse characteristics. After pre-training, the network weights are fixed, and the approximate PPD for a new training set can be computed through a single forward pass without additional training or tuning. The PPD quantifies uncertainty about the label $y$. However, it mixes the aleatoric and epistemic components of uncertainty (Hüllermeier and Waegeman,, 2021). From the PPD alone, it is impossible to disentangle these parts. Consequently, PFNs do not provide uncertainty estimates for predictive summaries, such as conditional mean, probabilities, and quantiles (see Figure 1).
@@ 2.2 Bayesian Inference
$\operatorname{PPD}(y|x)=\int p_{\theta}(y|x)\pi(\theta|\mathcal{D}_{n})\,d\theta.$
In contrast to the PPD, the posterior $\pi(\theta|\mathcal{D}_{n})$ also quantifies uncertainty for other interest quantities. For example, the posterior distribution for the conditional mean $\mu(x)=\int y\,p_{\theta}(y|x)dy$ is given by
$\Pi(\mu(x)\in A)=\int\mathds{1}\left\{\int y\,p_{\theta}(y|x)dy\in A\right\}\pi(\theta|\mathcal{D}_{n})\,\,d\theta,$
@@ 2.3 Martingale Posteriors
Martingale posteriors were recently introduced by Fong et al., (2023) as a new method for Bayesian uncertainty quantification. Its core idea is to reverse the direction of the Bayesian inference. In classical Bayesian inference, the posterior is derived from a prior and likelihood, which then implicitly leads to the PPD. MP inference starts from the PPD and leaves the prior $\pi(\theta)$ implicit. An appropriate sampling scheme and Doob’s theorem then allow us to derive posteriors for virtually all quantities of interest (e.g., the conditional mean $\mu(x)$).
@@
$y_{n+1}\sim p(\cdot|y_{1:n}),\quad y_{n+2}\sim p(\cdot|y_{1:(n+1)}),$ (1)
$y_{n+3}\sim p(\cdot|y_{1:(n+2)}),\quad\dots$ (2)
NN times, which yields a sample $y_{(n+1):(n+N)}$ drawn from the predictive joint distribution
@@
$\Pi(\theta\in A|\mathcal{D}_{n})=\int\mathds{1}\{\theta(F_{\infty})\}d\Pi(F_{\infty}|\mathcal{D}_{n}),$
where $A$ is any Borel set on the space where the parameter $\theta$ lives. Furthermore, Doob’s theorem (Doob,, 1949) implies that $\Pi(\theta|\mathcal{D}_{n})$ coincides with the classical Bayes posterior for the prior $\pi(\theta)$ implicit in the PPD (Fong et al.,, 2023).
@@ 3 EFFICIENT MARTINGALE POSTERIORS FOR PRIOR-DATA FITTED NETWORKS
Martingale posteriors allow for Bayesian inference directly from the PPD. PFNs approximate the PPD, so using PFNs to construct a martingale posterior seems natural. However, there are two problems. First, modern PFNs are based on transformer architectures that require $\Omega(n^{2})$ operations for a forward pass on a training set size of $n$. Iteratively computing $p(y|y_{1:(n+k)})$ for $k=1,\dots,N$ thus has complexity $\Omega(N^{3})$, which is prohibitive. Second, Falck et al., (2024) found that modern transformer-based LLMs substantially deviate from the martingale property
$\mathbb{E}[p(y|y_{1:(n+k)})|y_{1:n}]=p(y|y_{1:n}).$
Also TabPFN fails in this regard.
@@
LABEL:{fig:martprop} (experiment description in Section 4.1) shows a QQ-plot of the average CDF resulting from PFN’s PPD at iteration 0 vs. the CDF after $k$ forward iterations of the sampling procedure. The martingale property is satisfied when this distribution remains constant, i.e., all lines lie on the $x=y$ diagonal. However, we see an apparent change in the distribution over the different sampling horizons. This reinforces that PFNs approximate Bayesian inference only in a weak sense, even though they were conceived from Bayesian ideas (Hollmann et al.,, 2023; Müller et al.,, 2021).
Without the martingale property, the MP sampling procedure leads to meaningless results. Instead, we propose to use the PPD implied by the PFN only as a starting point for the sampling scheme. This PPD is then iteratively updated using the nonparametric Gaussian copula approach of Fong et al., (2023), which ensures the martingale property. The resulting martingale posterior effectively treats the PFN output as an informed prior, without requiring the PFN to provide coherent posterior updates.
@@ 3.1 Martingale Posteriors for Conditional Inference
We extend the unconditional sampling scheme outlined in the previous section to the conditional inference setting. Fong et al., (2023) already proposed one such extension. Their scheme involves forward sampling of the features $x_{(n+1):(n+N)}$. The distribution of the features isn’t of primary interest but significantly complicates the sampling procedure, which the authors resolved through heuristic simplifications. In contrast to Fong et al., (2023), we propose to sample only the labels $y_{(n+1):(n+N)}$ conditional on the event that $x_{n+k}=x$, for a fixed value of $x$ and all $k=1,\dots,N$.
@@
$\Pi(\theta(x)\in A|\mathcal{D}_{n})=\int\mathds{1}\{\theta(F_{\infty,x})\}d\Pi(F_{\infty,x}|\mathcal{D}_{n}).$
Common examples of the parameter $\theta(x)$ are the conditional mean $\theta(x)=\int y\,dP(y|x)dy$ or a conditional $\varrho$-quantile $\theta(x)=P^{-1}(\varrho|x)$.
@@ 3.2 Efficient PPD Updates Based on the Gaussian Copula
Observe that $p_{0}(y)=p(y|x,\mathcal{D}_{n})$ is the PPD approximated by the PFN $g_{\theta}$. However, the following update distributions $p_{1},p_{2},\dots$ are generally intractable. To alleviate this, Fong et al., (2023) proposed a computationally efficient, nonparametric method based on Dirichlet Process Mixture Models (DPMMs) and a copula decomposition of the conditional $p_{k}$. Specifically, we set
$P_{k}(y)=$ $(1-\alpha_{n+k})P_{k-1}(y)$ (3)
$+\alpha_{n+k}H_{\rho}(P_{k-1}(y),P_{k-1}(y_{n+k})),$
where $P_{k}$ is the CDF corresponding to $p_{k}$, $\alpha_{i}=\left(2-{i}^{-1}\right)(i+1)^{-1}$,
$H_{\rho}(u,v)=\Phi\left(\frac{\Phi^{-1}(u)-\rho\Phi^{-1}(v)}{\sqrt{1-\rho^{2}}}\right),$
and $\Phi$ is the standard normal cumulative distribution function. The method has a hyperparameter $\rho$, corresponding to a bandwidth that smoothes the updates. Fong et al., (2023) proposed to tune this by maximizing the likelihood of the updated densities over the observed data. To increase the alignment of the updates with the PFN baseline, we simulate a new sample from the PFN and optimize the bandwidth on the simulated data. More details can be found in Section B.1
3.3 Flexible Learning Rate Schedules
Fong et al., (2023) suggested the learning rate $\alpha_{i}=(2-i^{-1})(i+1)^{-1}$ for the unconditional context, where distribution functions can be estimated with convergence rate $n^{-1/2}$. As we shall see in Proposition 3 below, the learning rate $\alpha_{i}$ determines the contraction rate of the posterior. Our setting, however, is conditional on $x\in\mathbb{R}^{d}$, where nonparametric methods such as TabPFN must converge slower than $n^{-1/2}$, with rates depending on the dimension and smoothness of the estimation problem (see, e.g., Stone,, 1982).
To accommodate such settings, we allow for more flexible learning rate schedules that can be calibrated to match the behavior of the PFN. Specifically, we parametrize the learning rate as
$\alpha_{n+k}=C(n+k+1)^{-\beta},\quad C\in(0,\infty),\beta\in(0,1].$
where $C$ and $\beta$ should be chosen such that the behavior of the update (3) roughly matches the behavior of the PFN. In our experiments, we implement the following three schedules
  • default: $C=2$, $\beta=1$ as in (Fong et al.,, 2023),
  • type 1: $C=10^{-(d-1)/(d+4)}$, $\beta=1/2+d/(d+4)$,
  • type 2: $C=10^{-2d/(d+4)}$, $\beta=d/(d+4)$.
The constants $C$ are set to normalize the schedules to a comparable starting scales. We refer to our discussion after Proposition 3 in the next section for a motivation of the exponent $\beta$, and to Section 4.4 for empirical evaluation.
@@ 3.4 Theoretical Properties
Despite the simplicity of the updates given in (3), they provide essential theoretical guarantees. The following is a direct consequence of Theorem 3 by Fong et al., (2023).
Proposition 1.
@@
By de Finetti’s theorem (e.g., Theorem 1.49 in Schervish,, 2012), it then follows that there is a random variable $\Theta$ such that $(z_{1},z_{2},\dots)$ is conditionally iid given $\Theta$. The distribution of $\Theta$ can be interpreted as an implicit prior. In our setting, this prior depends both on the initial PPD implied by the PFN and the Gaussian copula updates specified by (3). In particular, the following result follows from Proposition 1 above and Theorem 2.2 of Berti et al., (2004).
Proposition 2.
@@
Proposition 3.
For any $y\in\mathbb{R}$, $N\geq 0$, the following holds with probability at least $1-\delta$:
$|P_{\infty,x}(y)-P_{N}(y)|\leq C\sqrt{\frac{\log(2/\delta)}{2\beta-1}}(n+N)^{-\beta+1/2}.$
The proof, given in Appendix A, is based on a time-uniform version of the Azuma-Hoeffding concentration inequality (Howard et al.,, 2020), the martingale property of the copula updates, and the learning rate schedule. Importantly, time-uniform martingale concentration allows us to directly bound the deviation of interest $|P_{\infty,x}-P_{N}|$ as opposed to $|P_{M}-P_{N}|$ with $M\geq N$ as in Fong et al., (2023).
Proposition 3 quantifies how well the distribution $P_{N}$ approximates $P_{\infty,x}$ after $N$ updates. Setting $N=0$ corresponds to the case where $P_{N}$ equals the initial PPD given only the observed data. In our setting, this is the output from the PFN. The random limit $P_{\infty,x}$ fluctuates around this point estimate, and the spread of these fluctuations characterizes the spread of the martingale posterior for $\theta(P(\cdot|x))$.
3.5 The Role of $\beta$
The case $\beta=1$ corresponds to the choice $\alpha_{i}\sim i^{-1}$ proposed by Fong et al., (2023) in the unconditional context with contraction rate of $n^{-1/2}$. Contraction rates for nonparametric regression problems are typically much slower than $n^{-1/2}$ (Ghosal and Van der Vaart,, 2017, Chapter 9), however. To accommodate such slow contraction rates, it is therefore crucial to allow for learning rates with $\beta<1$. The choice $\beta=d/(d+4)$ of our type 1 schedule corresponds to the optimal contraction rate we can expect in a nonparametric regression problem with $d$ features and twice continuously differentiable regression curve. However, Nagler, (2023) demonstrated that current PFN architectures yield inconsistent estimators, as their bias does not vanish as $n\to\infty$. This suggests that to get valid inferences, the posterior should never contract fully. This happens in the regime $\beta\in(0,1/2]$, where the bound in Proposition 3 becomes vacuous, but the algorithm still converges to a random limit $P_{\infty,x}$ because of Proposition 2. This motivates the choice $\beta=4/(d+4)$, which has $\beta\leq 1/2$ for the regime $d\geq 4$, where the bias becomes notable already on moderate sample sizes.
@@ 3.6 Computation
In practice, we can only sample finite sequences and replace the MP by its finite approximation. The procedure is summarized in Algorithm 1. The overall run-time complexity of the algorithm is $O(BN)$, where $B$ is the number of replications, and $N$ is the length of one ‘chain’. Based on the considerations in the previous subsection, we recommend generating around $N=200$ forward samples in each replication. The computations are independent across the outer loop iterations ($b=1,\dots,B$) and straightforward to parallelize; the inner loop ($k=1,\dots,N$) must be run sequentially. Except for the initial computation of $\widehat{\operatorname{PPD}}(y\mid x)$, the runtime is independent of the dataset size $n$.
@@
8:  Compute $\widehat{P}_{N}^{(b)}(y)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\Bigl\{y_{n+i}^{(b)}\leq y\Bigr\}$.
@@
[…]